# Hitting probability for Reflected Brownian Motion at Small Target

### v2

@@ 7. Appendix: Proofs of technical details
Notice that for $X^{x}_{t}$ inside domain $\Omega$, it satisfies Stochastic differential equation $\mathrm{d}X^{x}_{t}=\mathrm{d}W_{t}$. Write $\Psi_{m}=f=(f^{1},f^{2},...,f^{n})$ and $\Phi_{k}=g=(g^{1},...,g^{n})$, $W_{t}=(W^{1}_{t},...,W^{n}_{t})$ where $W^{i}_{t}$ are independent standard 1 dimensional Brownian motion. Apply Itô’s formua,
@@
where $\delta_{1i}=0$ for $i\neq 1$ and $1$ for $i=1$. We abbreviate starting position to make our equation clean. We also abbreviate
@@
where $\left(f^{-1}\right)^{p}$ is interpreted as the $p^{\mathrm{th}}$ component of $f^{-1}$. Notice that by written in multiplication form of Jacobian matrix the first term on RHS is just $\sum^{n}_{k=1}\frac{\partial g^{i}}{\partial x_{k}}\mathrm{d}W^{k}_{t}$ and second term is $\frac{1}{2}\Delta g^{i}\mathrm{d}t$. We also note that
@@
so $f^{-1}_{j}$ satisfies a Lipchitz condition. As we only have 4 $j$ we have the uniform Lipchitz condition. Also from here all partial derivative are nicely bounded by total derivative of $f^{-1}_{j}$. By Cauchy-Riemann Equation we see $J_{j}J_{j}^{T}=c\mathbb{I}$ where $c=|(f^{-1}_{j})^{\prime}|^{2}>0$, $J_{j}$ is Jacobian of $f^{-1}_{j}$ and $\mathbb{I}$ is the $2\times 2$ identity matrix. Also $f^{-1}_{j}$ are all harmonic, so property 4 is satisfied. Thus $\mathbb{D}$ is a SULD. For the case of general $\Omega$ in Definition 2.8, by Riemann mapping theorem we may choose a conformal bijection $f$ maps $\mathbb{D}$ onto $\Omega$. Now we introduce a lemma
@@
Together by same reasons listed in about proof that $\mathbb{D}$ satisfies property 4, we see property 4 is satisfied and thus $\Omega$ is a SULD. ∎
@@
Suppose $\Omega$ and $\Omega^{\prime}$ are two simply connected open subset of $\mathbb{H}$ and $B^{z_{0}}_{t}$ is the RBM in $\mathbb{H}$ where $z_{0}\in\Omega$ with the associated local time process $\xi^{z_{0}}_{t}$. Let $f$ be a conformal automorphism of $\mathbb{H}$. Let $\tau_{\Omega}=\inf\{t\geq 0:B^{z_{0}}_{t}\in\partial\Omega\}$. Then with probability one, $(f(B^{z_{0}}_{t}),t\in[0,\tau))$ is equal to the time-changed RBM in $\mathbb{H}$.
@@
with zero covariation. Let
@@
8. Acknowledgement
I would like to express my deepest gratitude to prof. Dmitry Belyaev as my supervisor of this summer project. He advised me the topic of this project and supervised me throughout the project.

### v4

@@ 7. Appendix: Proofs of technical details
Notice that for $X^{x}_{t}$ inside domain $\Omega$, it satisfies Stochastic differential equation $\mathrm{d}X^{x}_{t}=\mathrm{d}W_{t}$. Write $\Psi_{m}=f=(f^{1},f^{2},...,f^{n})$ and $\Phi_{k}=g=(g^{1},...,g^{n})$, $W_{t}=(W^{1}_{t},...,W^{n}_{t})$ where $W^{i}_{t}$ are independent standard 1 dimensional Brownian motion. Apply Itô’s formula,
@@
where $\delta_{1i}=0$ for $i\neq 1$ and $1$ for $i=1$. We abbreviate the starting position to make our equation clean. We also abbreviate
@@
where $\left(f^{-1}\right)^{p}$ is interpreted as the $p^{\mathrm{th}}$ component of $f^{-1}$. Notice that by written in multiplication form of the Jacobian matrix, the first term on RHS is just $\sum^{n}_{k=1}\frac{\partial g^{i}}{\partial x_{k}}\mathrm{d}W^{k}_{t}$ and the second term is $\frac{1}{2}\Delta g^{i}\mathrm{d}t$. We also note that
@@
so $f^{-1}_{j}$ satisfies a Lipchitz condition. As we only have 4 $j$ we have the uniform Lipchitz condition. Also, from here, all partial derivatives are nicely bounded by the total derivative of $f^{-1}_{j}$. By Cauchy-Riemann Equation we see $J_{j}J_{j}^{T}=c\mathbb{I}$ where $c=|(f^{-1}_{j})^{\prime}|^{2}>0$, $J_{j}$ is Jacobian of $f^{-1}_{j}$ and $\mathbb{I}$ is the $2\times 2$ identity matrix. Also, $f^{-1}_{j}$ are all harmonic, so property 4 is satisfied. Thus $\mathbb{D}$ is a SULD. For the case of general $\Omega$ in Definition 2.8, by Riemann mapping theorem, we may choose a conformal bijection $f$ maps $\mathbb{D}$ onto $\Omega$. Now we introduce a lemma
@@
Together by the same reasons listed in about proof that $\mathbb{D}$ satisfies property 4, we see property 4 is satisfied, and thus $\Omega$ is a SULD. ∎
@@
Suppose $\Omega$ and $\Omega^{\prime}$ are two simply connected open subset of $\mathbb{H}$ and $B^{z_{0}}_{t}$ is the RBM in $\mathbb{H}$ where $z_{0}\in\Omega$ with the associated local time process $\xi^{z_{0}}_{t}$. Let $f$ be a conformal automorphism of $\mathbb{H}$. Then the process $f(B^{z_{0}}_{t})$ is indistinguishable from a time-changed RBM in $\mathbb{H}$.
@@
with zero quadratic covariation. Let
@@
8. Acknowledgment
I want to express my deepest gratitude to Prof. Dmitry Belyaev as my supervisor of this summer project. He advised me on the topic of this project and supervised me throughout the project.